# RIS-Enabled Spoofing Against Adversary Sensing: CRB-Maximizing Design and Decoying Analysis

### v1

@@ Abstract
This paper studies the capability of a Reconfigurable Intelligent Surface (RIS), when transparently covering a User Equipment (UE), to deceive an adversary monostatic radar system. A compact RIS kernel model that explicitly links the radar’s angular response to the RIS phase profile is introduced, enabling an analytical investigation of the Angle of Arrival (AoA) estimation accuracy with respect to the kernel’s power. This model is also leveraged to formulate an RIS-based spoofing design with the dual objective to enforce strict nulls around the UE’s true reflection AoA and maximize the channel gain towards a decoy direction. The RIS’s deception capability is quantified using point-wise and angle-range robust criteria, and a configuration-independent placement score guiding decoy selection is proposed. Selected numerical results confirm deep nulls at the true reflection AoA together with a pronounced decoy peak, rendering maximum-likelihood sensing at the adversary radar unreliable.
@@ I Introduction
Reconfigurable Intelligent Surfaces (RISs) are engineered metasurfaces whose constituent subwavelength-space elements impose programmable phase shifts on incident waves [1], enabling, when placed accordingly [2], a wide variety of performance improvements, range channel shaping [3] and massive user access [4] to seamless non Line-of-Sight (LoS) localization and sensing [5]. Within the European Telecommunications Standards Institute (ETSI), a dedicated industry specification group [6] is currently issuing detailed reports on use cases, single- and multi-functional architectures, orchestration protocols, and field trials, paving the way for the technology’s consideration in the future 3GPP releases [7].
In the context of the emerging paradigm of Integrated Sensing And Communications (ISAC) [8], and, particularly, in the respective concerns for sensing security and privacy [9], RISs have been lately deployed for sensing protection against adversary radars. In [10], RIS-aided stealth was studied with the goal to minimize the echo power at hostile Receivers (RXs), presenting, under idealized LoS assumptions, closed-form designs for a single monostatic radar as well as low-complexity solvers based on the mean square error criterion for multi-radar scenarios. The authors in [11] utilized an RIS to suppress returns along the true direction from a malicious monostatic radar while redirecting energy towards clutter for the radar to infer a false Angle-of-Arrival (AoA). A hybrid RIS designed was deployed in [12] to shield a region from an adversary while preserving communications and legitimate sensing. In particular, an alternating optimization strategy to jointly configure the transmit and RIS-induced beamforming was presented. Very recently, in [13], an RIS was optimized to occultate user positions from wiretappers, while maintaining both sensing and desired communication performance between the user and a legitimate Base Station (BS). However, the state of the art either relies on clutter to mislead an adversary from sensing or suppresses echo power, while there is a lack of a Cramér-Rao Bound (CRB) analysis quantifying the fundamental deception limits of RIS-based spoofing.
@@ III Proposed RIS-Enabled Spoofing Design
In this section, we first derive the CRB for the AoA estimation of the reflected signal from the RIS-coated UE at the adversary monostatic radar BS. Inspired by this bound’s scaling behavior, a novel optimization framework that jointly nulls the backscattered signal towards the BS to conceal the UE and synthesizes a decoy echo at a prescribed AoA to mislead the respective estimation is presented.
@@ III-B Problem Formulation and Solution
This paper’s spoofing objective is to design the RIS so as to enforce an erroneous UE position estimation at the radar BS. In particular, the goal is to null the radar’s response at the true AoA $\theta_{\text{true}}$ and direct the incident radar wave towards a decoy angle $\theta_{\text{fake}}$. Using (4), we define the decoy kernel vector $\mathbf{w}\triangleq\bm{\upsilon}(\theta_{\text{fake}},\theta_{\text{true}})$ as well as the vector with the nulling angles $\bm{\Theta}_{\text{null}}\triangleq\{\theta_{1},\ldots,\theta_{K}\}$, where $\theta_{k}\in[\theta_{\text{true}}-\Delta,\theta_{\text{true}}+\Delta]$ $\forall k=1,\ldots,K$ with $\Delta$ being a width that specifies the RIS-centered angle nulling window over which we will enforce $\bar{\beta}(\theta_{k};\bm{\omega})\approx 0$ $\forall k$. We next collect the RIS kernel vectors associated with the angle nulling window [^2] We impose a null over an angular range around $\theta_{\text{true}}$ rather than at a single point because point nulls can be inaccurate. Even with accurate AoA knowledge at the RIS, the existence of noise can shift the radar’s response by a small offset. A banded, robust null ensures low radar response, and thus UE AoA concealment, throughout a neighborhood of $\theta_{\text{true}}$ [17]. as follows:
@@
The projection onto $\mathcal{T}$ is achieved by the element-wise phase normalization operation: $\bm{\Pi}_{\mathcal{T}}(\mathbf{x})=e^{\jmath\angle(\mathbf{x})}$. A simple scheme solving $\mathcal{P}_{3}$ alternates between the subspace projection onto $\mathcal{S}$ and the unit-modulus projection onto $\mathcal{T}$ [19, 20]. This procedure is summarized in Algorithm 1, where $I_{\text{max}}$ is the maximum number of iterations, $\varepsilon_{\text{null}}$ is the nulling tolerance, and $\gamma\in(0,1)$ is a relaxation parameter that, at each step, nudges the iterate towards the decoy vector.
@@ V-A Position Estimation Performance
Figure 1 contrasts the RIS beampattern gain and the corresponding ML-based AoA spectrum before and after the proposed RIS-aided spoofing. As seen, with no phase control (i.e., $\bm{\omega}]_{m}=1\forall m$), the beampattern returns energy towards the true UE direction $\theta_{\text{true}}$. However, after the proposed RIS optimization, energy is redirected to the prescribed decoy AoA while a deep, banded null is carved around $\theta_{\text{true}}$. The ML spectrum mirrors this redistribution: it is flat and uninformative near $\theta_{\text{true}}$ exhibiting a sharp peak at the decoy AoA, systematically pulling the estimator away from the actual target. This behavior aligns with the FI/CRB analysis: suppressing kernel power at $\theta_{\text{true}}$ collapses local FI, inflating the CRB, whereas concentrating power at the decoy tightens the bound and stabilizes the ML peak. The banded null enhances robustness to small angular mismatches, and subdued sidelobes limit leakage that might otherwise reveal $\theta_{\text{true}}$.
@@ V-B Investigation of the $\rho$-Deception Criterion
Figure 3 illustrates the trade-off between $\theta_{\text{fake}}$’s amplitude and residual leakage by plotting the leakage ratio $L_{\rm true}(\bm{\omega})/|\bar{\beta}(\theta_{\text{fake}};\bm{\omega})|$ versus the decoy angle, together with band-robust $\rho$-deception thresholds $\sqrt{\kappa(\theta_{\text{fake}})/(\rho\,\kappa_{\min})}$ for the optimized RIS-based spoofing design of Fig. 1. Decoy angles where the curve falls below a threshold achieve the corresponding $\rho$-deception level; those above fail it. t is shown that the optimized RIS profile attains a low ratio at the chosen decoy (here $\theta_{\text{fake}}=-48^{\circ}$), evidencing effective deception, yet it is not the global minimum; this reflects the priority of enforcing a hard, banded null around $\theta_{\text{true}}=20^{\circ}$, which necessarily limits decoy gain and elevates worst-case leakage. Consistent with Fig. 2, the beam towards the decoy position is relatively broad. It can, thus, be concluded from Fig. 3 that the proposed spoofing design provides deeper, band-robust nulls enhancing concealment around $\theta_{\text{true}}$, but it limits the achievable decoy gain, with the realized strength governed by $\mathcal{P}_{3}$’s solution efficiency.
@@
Finally, the RIS-agnostic upper bound $\rho_{\mathrm{UB}}\left(\theta;\overline{L}\right)$ in (18) is evaluated with respect to the decoy angle $\theta_{\text{fake}}$ in Fig. 4. It can be seen that the peak locations are governed by $\Phi(\theta)$: they occur when the RIS can radiate strongly after projection onto the true AoA null subspace (i.e., large $\eta(\theta)$) and the radar BS array/illumination factor is highly directive (i.e., large $\kappa(\theta)$). As observed, varying $\overline{L}$, merely rescales the curve without shifting these maxima: the same angles remain optimal targets for deception irrespective of the leakage budget, while the pronounced through across the nulled band around $\theta_{\text{true}}$ confirms that robust nulling renders deception unattainable there.

### v2

@@ Abstract
This paper studies the capability of a Reconfigurable Intelligent Surface (RIS), when transparently covering a User Equipment (UE), to deceive an adversary monostatic radar system. A compact RIS kernel model that explicitly links the radar’s angular response to the RIS phase profile is introduced, which enables an analytical investigation of the Angle of Arrival (AoA) estimation accuracy with respect to the kernel’s power. This model is also leveraged to formulate an RIS-based spoofing design with the dual objective to enforce strict nulls around the UE’s true reflection AoA as well as maximize the channel gain toward a decoy direction. In addition, the RIS’s deception capability is quantified using point-wise and angle-range robust criteria, and a configuration-independent placement score guiding decoy selection is proposed. The presented numerical results confirm deep nulls at the true reflection AoA together with a pronounced decoy peak, a fact that eventually renders maximum-likelihood sensing at the adversary radar unreliable.
@@ I Introduction
Reconfigurable Intelligent Surfaces (RISs) are engineered metasurfaces whose constituent subwavelength-space elements impose programmable phase shifts on incident waves [1, 2], enabling, when placed accordingly [3], a wide variety of performance improvements, range channel shaping [4] and massive user access [5] to seamless non Line-of-Sight (LoS) localization and sensing [6]. Within the European Telecommunications Standards Institute (ETSI), a dedicated industry specification group [7] is currently issuing detailed reports on use cases, single- and multi-functional architectures, orchestration protocols, and field trials, paving the way for the technology’s consideration in the future 3GPP releases [8].
In the context of the emerging paradigm of Integrated Sensing And Communications (ISAC) [9], and, particularly, in the respective concerns for sensing security and privacy [10], RISs have been lately deployed for sensing protection against adversary radars. In [11], RIS-aided stealth was studied with the goal to minimize the echo power at hostile Receivers (RXs), presenting, under idealized LoS assumptions, closed-form designs for a single monostatic radar as well as low-complexity solvers based on the mean square error criterion for multi-radar scenarios. The authors in [12] utilized a RIS to suppress returns along the true direction from a malicious monostatic radar while redirecting energy toward clutter for the radar to infer a false Angle-of-Arrival (AoA). A hybrid RIS designed was deployed in [13] to shield a region from an adversary while preserving communications and legitimate sensing. In particular, an alternating optimization strategy to jointly configure the transmit and RIS-induced beamforming was presented. Very recently, in [14], an RIS was optimized to occultate user positions from wiretappers, while maintaining both sensing and desired communication performance between the user and a legitimate Base Station (BS). However, the state of the art either relies on clutter to mislead an adversary from sensing or suppresses echo power, while there is a lack of a Cramér-Rao Bound (CRB) analysis quantifying the fundamental deception limits of RIS-based spoofing.
@@ III Proposed RIS-Enabled Spoofing Design
In this section, we first derive the CRB for the AoA estimation of the reflected signal from the RIS-coated UE at the adversary monostatic radar BS. Inspired by this bound’s scaling behavior, a novel optimization framework that jointly nulls the backscattered signal toward the BS to conceal the UE and synthesizes a decoy echo at a prescribed AoA to mislead the respective estimation is presented.
@@ III-B Problem Formulation and Solution
This paper’s spoofing objective is to design the RIS so as to enforce an erroneous UE position estimation at the radar BS. In particular, the goal is to null the radar’s response at the true AoA $\theta_{\text{true}}$ and direct the incident radar wave toward a decoy angle $\theta_{\text{fake}}$. Using (4), we define the decoy kernel vector $\mathbf{w}\triangleq\bm{\upsilon}(\theta_{\text{fake}},\theta_{\text{true}})$ as well as the vector with the nulling angles $\bm{\Theta}_{\text{null}}\triangleq\{\theta_{1},\ldots,\theta_{K}\}$, where $\theta_{k}\in[\theta_{\text{true}}-\Delta,\theta_{\text{true}}+\Delta]$ $\forall k=1,\ldots,K$ with $\Delta$ being a width that specifies the RIS-centered angle nulling window over which we will enforce $\bar{\beta}(\theta_{k};\bm{\omega})\approx 0$ $\forall k$. We next collect the RIS kernel vectors associated with the angle nulling window [^2] We impose a null over an angular range around $\theta_{\text{true}}$ rather than at a single point because point nulls can be inaccurate. Even with accurate AoA knowledge at the RIS, the existence of noise can shift the radar’s response by a small offset. A banded, robust null ensures low radar response, and thus UE AoA concealment, throughout a neighborhood of $\theta_{\text{true}}$ [18]. as follows:
@@
The projection onto $\mathcal{T}$ is achieved by the element-wise phase normalization operation: $\bm{\Pi}_{\mathcal{T}}(\mathbf{x})=e^{\jmath\angle(\mathbf{x})}$. A simple scheme solving $\mathcal{P}_{3}$ alternates between the subspace projection onto $\mathcal{S}$ and the unit-modulus projection onto $\mathcal{T}$ [20, 21]. This procedure is summarized in Algorithm 1, where $I_{\text{max}}$ is the maximum number of iterations, $\varepsilon_{\text{null}}$ is the nulling tolerance, and $\gamma\in(0,1)$ is a relaxation parameter that, at each step, nudges the iterate toward the decoy vector.
@@ V-A Position Estimation Performance
Figure 1 contrasts the RIS beampattern gain and the corresponding ML-based AoA spectrum before and after the proposed RIS-aided spoofing. As seen, with no phase control (i.e., $\bm{\omega}]_{m}=1\forall m$), the beampattern returns energy toward the true UE direction $\theta_{\text{true}}$. However, after the proposed RIS optimization, energy is redirected to the prescribed decoy AoA while a deep, banded null is carved around $\theta_{\text{true}}$. The ML spectrum mirrors this redistribution: it is flat and uninformative near $\theta_{\text{true}}$ exhibiting a sharp peak at the decoy AoA, systematically pulling the estimator away from the actual target. This behavior aligns with the FI/CRB analysis: suppressing kernel power at $\theta_{\text{true}}$ collapses local FI, inflating the CRB, whereas concentrating power at the decoy tightens the bound and stabilizes the ML peak. The banded null enhances robustness to small angular mismatches, and subdued sidelobes limit leakage that might otherwise reveal $\theta_{\text{true}}$.
@@ V-B Investigation of the $\rho$-Deception Criterion
Figure 3 illustrates the trade-off between $\theta_{\text{fake}}$’s amplitude and residual leakage, by plotting the leakage ratio $L_{\rm true}(\bm{\omega})/|\bar{\beta}(\theta_{\text{fake}};\bm{\omega})|$ versus the decoy angle, together with band-robust $\rho$-deception thresholds $\sqrt{\kappa(\theta_{\text{fake}})/(\rho\,\kappa_{\min})}$ for the optimized RIS-based spoofing design of Fig. 1. Decoy angles where the curve falls below a threshold achieve the corresponding $\rho$-deception level; those above fail it. It is also shown that the optimized RIS profile attains a low ratio at the chosen decoy (here $\theta_{\text{fake}}=-48^{\circ}$), evidencing effective deception, yet it is not the global minimum. This behavior reflects the priority of enforcing a hard, banded null around $\theta_{\text{true}}=20^{\circ}$, which necessarily limits decoy gain and elevates worst-case leakage. Consistent with Fig. 2, the beam toward the decoy position is relatively broad. It can, thus, be concluded from Fig. 3 that the proposed spoofing design provides deeper, band-robust nulls enhancing concealment around $\theta_{\text{true}}$, but it limits the achievable decoy gain, with the realized strength governed by the efficiency of the solution of the $\mathcal{P}_{3}$ problem.
@@
Finally, the RIS-agnostic upper bound $\rho_{\mathrm{UB}}\left(\theta;\overline{L}\right)$ in (18) is evaluated with respect to the decoy angle $\theta_{\text{fake}}$ in Fig. 4. It can be seen that the peak locations are governed by $\Phi(\theta)$: they occur when the RIS can radiate strongly after projection onto the true AoA null subspace (i.e., large $\eta(\theta)$) and the radar BS array/illumination factor is highly directive (i.e., large $\kappa(\theta)$). As observed, varying $\overline{L}$, merely rescales the curve without shifting these maxima: the same angles remain optimal targets for deception irrespective of the leakage budget, while the pronounced dip across the nulled band around $\theta_{\text{true}}$ confirms that robust nulling renders deception unattainable there.